Use the given information to find the indicated probability. A and B are mutually exclusive. P(A) = .1, P(B) = .7. Find P((A ∪ B)').
Answer #1

Solution :

If A and B are mutually exclusive then ,

P(A \cup B) = P(A) + P(B)

P(A \cup B)c = 1 - P(A \cup B) = 1 - P(A) - P(B)

= 1 - 0.1 - 0.7

= 0.2